OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Find the inverse of f(x) = 3x + 4. Is the inverse a function? Explain

jimthompson5910 (jim_thompson5910):

f(x) = 3x + 4 y = 3x + 4 ... replace f(x) with y x = 3y + 4 ... swap x and y Now solve for y

OpenStudy (abbycross167):

@jim_thompson5910 y=−4/3+x/3

jimthompson5910 (jim_thompson5910):

yes or y = (x-4)/3

jimthompson5910 (jim_thompson5910):

to answer the question `Is the inverse a function?` graph the inverse and see if it passes the vertical line test